THE ULTRAVIOLET SPECTRUM OF THE FeF RADICAL

\underline{STEPHEN M. KERMODE} and JOHN M. BROWN, The Physical and Theoretical Chemistry Laboratory, South Parks Road, Oxford, OX1 3QZ, United Kingdom.

Two band systems of the FeF radical have been recorded at Doppler-limited linewidths by Laser Induced Fluorescence. The (0,0) band of the ^{6}Π−X^{6}Δ was recorded at 323 nm. The system at 330 nm consisted of an overlap of the (1,0) band of the ^{6}Π−X^{6}Δ system and the (0,0) band of the ^{6}Φ−X^{6}Δ system. FeF was produced in a continuous flow by the reaction of fluorine atoms, produced by a microwave discharge of fluorine in helium, with iron pentacarbonyl. This reaction provides data for the lowest four sub-bands of each system, considerably extending our knowledge of these two electronic transitions compared with previous work by Pouilly et al.^{a}. The temperature of the sample in our experiments is slightly above ambient, about 450 K. The results have been analysed and combined with the millimeter-wave spectra^{b} for the ground state, ^{6}Δ, to produce better parameters for the excited states.

aB. Pouilly, J. Schamps, D. J. W. Lumley and R. F. Barrow, J. Phys B: Atom. Molec. Phys., 11, 2289, (1978).

bM. D. Allen and L. M. Ziurys, J. Chem. Phys., 106(9), 3494 (1997).